# Maximum principles for matrix-valued regular functions of a quaternionic variable

### v2

@@ Abstract.
A quaternionic matrix-valued regular function is a map $F:\Omega\rightarrow M_{n}(\mathbb{H})$ whose entries are regular functions of a quaternion variable, where $\Omega$ is a domain in $\mathbb{H}$. Our aim is to bring out some maximum norm principles for such functions. We derive a decomposition theorem for such functions and also prove a Caratheodory-Rudin type approximation theorem for functions in the quaternionic Schur class. This in turn yields that a $2\times 2$ norm one matrix-valued function can be approximated by quaternionic rational inner functions.
@@
Quaternions; symmetric slice domain; matrix-valued regular functions; complex adjoint matrix; maximizing vector; maximum principle; rational inner functions
@@ 2. Definitions and Preliminaries
We shall use the above notation $\mathbb{S}$ throughout this paper to denote only the $2$-sphere.
For $I\in\mathbb{S}$, we denote by $L_{I}$ the set $\mathbb{R}+\mathbb{R}I$, which we may identify with $\mathbb{C}$. We then define $\Omega_{I}$ to be the intersection of any domain $\Omega$ with $L_{I}$ and $f|_{I}$ to be the restriction of $f:\Omega\rightarrow\mathbb{H}$ on $\Omega_{I}$. A function $f:\Omega\rightarrow\mathbb{H}$ is said to be regular on a domain $\Omega$ if for any $I\in\mathbb{S}$, the function $f|_{I}$ is complex holomorphic on $\Omega_{I}$. Every regular function on $\mathbb{B}$ admits a nice power series representation, as defined below.
@@
An analagous definition of right regular functions exist; however, we work only with left regular functions in this paper. All of the definitions, facts and theorems stated below can be found in [8] or [9]. With the above definitions, any left polynomial over quaternions is an example of a regular function. Another simple example of a regular function over the quaternions is $\frac{1}{q}$, where $0\neq q\in\mathbb{H}$. There are functions that are not regular too. We shall exhibit one in Example 3.13.
Remark: From now on, we use the phrase regular function only to mean that it is a left regular function, to avoid confusion, although the definitions and facts presented below also hold for right regular functions.
We now recall the Splitting Lemma (see for instance [8] or [9]), which helps us to restrict a regular function to $L_{I}$, and hence get complex holomorphic function.
@@
A brief description of the results obtained are as below. After a brief introduction, the necessary definitions and preliminaries are brought out. This is followed by the section on main results, which is subdivided into four subsections for ease of reading. The first of these concerns maximum principles for Frobenius and operator norms of matrix-valued functions. Maximizing vectors and a decomposition result for matrix-valued regular functions are brought out next. We then discuss maximum principles involving singular values of matrix-valued functions. The last subsection concerns approximating a regular matrix-valued function by rational inner functions. We bring out a Caratheodory-Rudin type approximation result of a regular function of a quaternion variable and use it to prove that a $2\times 2$ matrix-valued regular function of a quaternion variable can be approximated by matrix-valued rational inner functions.
@@ 3.1. The Frobenius and Operator norm case
As pointed out in in Theorem $2.12$ of [9], if we work with an arbitrary domain $\Omega$ in $\mathbb{H}$, then any regular function will have a power series expansion in each $\Sigma$-ball in $\Omega$ and in such a case, the proof of Theorem 3.3 can be modified appropriately.
@@ 3.2. Maximizing vectors and a decomposition theorem
Let $\Omega$ be a slice domain in $\mathbb{H}$ and $F:\Omega\rightarrow M_{n}(\mathbb{H})$ a regular matrix-valued function. If there exists a $q_{0}$ in $\Omega_{I}$ such that $\|F(q)\|\leq\|F(q_{0})\|$ for all $q\in\Omega$ and $x_{0}$ is the maximizing vector for $F(q_{0})$, then $F^{k}(q_{0})x_{0}=0$ for all $k\geq 1$.
@@
Let the assumptions stated above hold. Choose $R>0$ such that $\Sigma(q_{0},R)$ contained in $\Omega$ and
@@
where $F_{n}\in M_{n}(\mathbb{H})$. Let $x\in\mathbb{H}$. Since $\mathbb{H}$ is a left Hilbert space, we have
@@
Let $\delta_{I}(q_{0},R)\subseteq\Sigma(q_{0},R)$ be the disk in $L_{I}$ of radius $R$ and centered at $q_{0}$. Then for any $q\in\delta_{I}(q_{0},R)$, choose $r<R$ such that $q=q_{0}+re^{it}$. We then have,
@@
By the regular power series representation (see [9]), $\|F(q)\|\leq\|F(q_{0})\|=\|F_{0}\|$ for all $q\in\delta_{I}(q_{0},R)$; in particular, for the vector $x$ and $r<R$, it follows that
@@
Replacing $x$ by the maximizing vector $x_{0}$, we get
@@
This implies that $F_{k}x_{0}=0$ for all $k\geq 1$ and hence $F^{k}(z_{0})x_{0}=0$ for all $k\geq 1$ and hence $F(q)x_{0}=F_{0}x_{0}=F(q_{0})x_{0}$ for all $q\in\delta_{I}(q_{0},R)$. We finally conclude from the identity principle (Theorem 2.6) that $F(q)x_{0}$ is constant on $\Omega$, as required. ∎
@@
It should be pointed out that we have assumed that the point $q_{0}\in\Omega_{I}$ (the point where the maximum is attained for $|F(q)|$) some $I\in\mathbb{S}$ in the theorems proved in this section and the previous one. However, this is merely a technical issue and the proofs will go through verbatim, with appropriate modifications, if $q_{0}\in\Omega$. This is because for any $q\in\mathbb{H}$, we can choose $I,J\in\mathbb{S}$ such that $q\in L_{I}$ by taking $v=x_{1}i+x_{2}j+x_{3}k$ (the imaginary part of $q$) and by letting $I=\dfrac{v}{|v|}$ so that $q=x_{0}+|v|I\in L_{I}$.
@@ 3.4. Approximation of a quaternionic matrix-valued function
This is the last section of the paper. We prove as an application of Theorem 3.8, a Caratheodory-Rudin type approximation theorem. We begin with the following two definitions.
@@
Let $\Omega$ be subset of $\mathbb{H}$. A function $f:\Omega\rightarrow\mathbb{H}$ is said to be a rational function if $f(q)=h(q)^{-*}*g(q)$ where $h(q)$ and $g(q)$ are (left) polynomials with coefficients from quaternions such that $h(q)$ has no zeros in $\Omega$.
Definition 3.16.
A function $f:\mathbb{B}\rightarrow M_{n}(\mathbb{H})$ is called an inner function if the boundary values of the function on the unit ball are unitary matrices everywhere.
We are now in a position to prove a Caratheodory-Rudin type approximation result. The proof uses a Runge type approximation theorem over quaternions. Runge theorem for slice hyperholomorphic functions was recently studied in [5].
@@
Let $\mathbb{B}$ be the open unit ball in $\mathbb{H}$ and $f:\mathbb{B}\rightarrow\overline{\mathbb{B}}$ be a regular function on $\mathbb{B}$. Then $f$ can be approximated by rational inner functions.
@@
By the splitting lemma (Lemma 2.3), any regular function $f$ can be written as $f_{J}=g+hK$, where $J,K\in\mathbb{S},\ J\perp K$ and $g,h:\mathbb{B}_{J}\rightarrow\overline{\mathbb{B}_{J}}$ being holomorphic functions on $\mathbb{B}_{J}$. By the Caratheodory-Rudin theorem for the complex case (see [2] for details), $g$ and $h$ can be approximated by rational inner functions in field of complex numbers, say $r_{1}$ and $r_{2}$, such that for $\epsilon>0$, we have $|g(z)-r_{1}(z)|<\epsilon/4$ and $|h(z)-r_{2}(z)|<\epsilon/4$ for all $z=x+yJ\in\mathbb{B}_{J}$. We can extend $r(x+yJ):=r_{1}(x+yJ)+r_{2}(x+yJ)K$ to the entire unit ball $\mathbb{B}$ by extension, say, $r(z)$ for $I\in\mathbb{S}$ (see Proposition $1.3.2$ of [8] for details) for $I\in\mathbb{S}$. Then,
$ext(r(q))=r(x+yI)=\frac{1}{2}\left[r(x+yJ)+r(x-yJ)+IJ\left[r(x-yJ)-r(x+yJ)\right]\right].$
@@
$|f(q)-r(q)|$ $=\Bigg|\frac{1}{2}[f(x+yJ)+f(x-yJ)]+IJ\left[f(x-yJ)-f(x+yJ)\right]$
@@
$|f(q)-r(q)|$ $=\Bigg|\frac{1}{2}\Bigg[(1-IJ)\Big[g(x+yJ)-r_{1}(x+yJ)+(h(x+yJ)-r_{2}(x+yJ))K\Big]$
$+(1+IJ)\Big[g(x-yJ)-r_{1}(x-yJ)+(h(x-yJ)-r_{2}(x-yJ))K\Big]\Bigg]\Bigg|$
$\leq\Big|g(x+yJ)-r_{1}(x+yJ)\Big|+\Big|h(x+yJ)-r_{2}(x+yJ)\Big|$
$+\Big|g(x-yJ)-r_{1}(x-yJ)\Big|+\Big|h(x-yJ)-r_{2}(x-yJ)\Big|$
@@
We thus conclude that $f$ can be approximated by rational inner functions over $\mathbb{H}$, as desired. ∎
@@
As defined in [3], regular functions that satisfy the hypothesis of the above theorem are said to be in quaternionic Schur class (see Definition $1.1$ of [3]). We have thus proved that any function that belongs to the quaternionic Schur class can be approximated by rational inner functions over $\mathbb{H}$. As pointed out in [3], both right and left Schur classes over quaternions coincide. As a final remark, we wish to point out that the quaternionic Schur class was studied in detail by Alpay et al (see Theorem $3.3$ and its consequences of [1]) and Bolotnikov (see Section $4$ of [4]).
We end this paper by proving the following result using the maximum norm principle.
@@
Let $F:\mathbb{B}\rightarrow M_{2}(\mathbb{H})$ be a regular function such that $\|F(q)\|=1$. Then $F$ can be approximated by $M_{2}(\mathbb{H})$-valued rational inner functions.
@@
Since $\|F(q)\|=1$, Theorem 3.8 tells us that there exists $2\times 2$ constant unitary matrices $U$ and $V$ and a regular function $G:\mathbb{B}\rightarrow M_{1}(\mathbb{H})\cong\mathbb{H}$ such that
$F(q)=U\begin{bmatrix}1&0\\ 0&G(z)\end{bmatrix}V^{*}.$
The Theorem 3.17 will then imply that there is an approximation of $G$ by rational inner functions, and hence $F$ can also be approximated by matrix-valued rational inner functions. ∎

### v3

@@ Abstract.
A quaternionic matrix-valued regular function is a map $F:\Omega\rightarrow M_{n}(\mathbb{H})$ whose entries are (left) regular functions of a quaternion variable, where $\Omega$ is a domain in $\mathbb{H}$. Our aim is to bring out some maximum norm principles for such functions. We derive an SVD type decomposition theorem for such functions, using the notion of maximizing vectors. Some maximum principles for singular values of matrix-valued regular function are brought out next. We then proceed to prove a Fisher type approximation theorem for regular functions $f:\mathbb{B}\rightarrow\overline{\mathbb{B}}$ that are continuous on $\partial\mathbb{B}$, in terms of convex combinations of finite Blaschke products over $\mathbb{H}$ ($\mathbb{B}$ being the quaternionic unit ball). This in turn yields a Fisher type approximation theorem for an $n\times n$ matrix-valued regular function on the quaternionic unit ball, where each entry of the matrix satisfies the same condition as above.
@@
Quaternions; symmetric slice domain; matrix-valued regular functions; complex adjoint matrix; maximizing vector; maximum principle; finite Blaschke products; Caratheodory and Fisher type approximation theorems
@@ 2. Definitions and Preliminaries
We shall use the above notation $\mathbb{S}$ throughout this paper to denote only the $2$-sphere. Notice that $\mathbb{S}$ belongs to the hyperplane $\{x_{0}=0\}$ and is therefore a $2$-dimensional sphere in $\mathbb{H}$. If $I\in\mathbb{S}$, then it can be easily verified that $I^{2}=-1$. For $I\in\mathbb{S}$, we denote by $L_{I}$ (or $C_{I}$) the set $\mathbb{R}+\mathbb{R}I$, which we may identify with $\mathbb{C}$. We then define $\Omega_{I}$ to be the intersection of any domain $\Omega$ with $L_{I}$.
@@
Let $U$ be an open subset of $\mathbb{H}$ and $f:U\rightarrow\mathbb{H}$ be a real differentiable function. $f$ is said to be left slice regular or left slice hyperholomorphic or simply left regular, if for every $I\in\mathbb{S}$, its restriction $f_{I}$ to $C_{I}$ satisfies
$\overline{\partial}_{I}f(x+Iy)=\frac{1}{2}\Bigg(\frac{\partial}{\partial x}+I\frac{\partial}{\partial y}\Bigg)f_{I}(x+Iy)I=0.$
An analagous definition of right regular functions also exist. However, we work only with left regular functions in this paper. It can be verified that the collection of all left regular functions is a right linear space over $\mathbb{H}$. It turns out that the condition of slice regularity given in Definition 2.2 holds if for a fixed $I\in\mathbb{S}$ if and only if for every $J\perp I,\ J\in\mathbb{S}$, there exist complex-valued holomorphic functions $F,G:U\cap C_{I}\rightarrow C_{I}$ such that $f_{I}(z)=F(z)+G(z)J$ for all $z=x+Iy\in U\cap C_{I}$. This equivalence comes through the Splitting Lemma, stated below.
@@
With the above definitions, any left polynomial over quaternions is an example of a regular function. Uniformly convergent power series also satisfy the conditions of Definition 2.2. That the converse is true over balls centered at the origin is given in the following theorem.
Theorem 2.4.
Let $B(0,R)$ be the open ball centered at 0 of radius $R$ in $\mathbb{H}$. A function $f:B(0,R)\rightarrow\mathbb{H}$ is left slice regular if and only if $f$ admits a power series representation $\sum_{n=0}^{\infty}q^{n}a_{n}$ that converges uniformly on $B(0,R)$.
We thus have the following equivalent definition of left regular functions over balls centered at the origin.
Definition 2.5.
A simple example of a non-polynomial regular function over $\mathbb{H}$ is $\frac{1}{q}$, where $0\neq q\in\mathbb{H}$. There are functions that are not regular too. We shall exhibit one in Example 3.13.
Remark: From now on, we use the phrase regular function only to mean that it is a left regular function, to avoid confusion, although some of the definitions and facts presented below also hold for right regular functions. All of the definitions, facts and theorems stated below can be found in [9] or [11].
@@
A brief description of the results obtained are as below. After a brief introduction, the necessary definitions and preliminaries are brought out. This is followed by the section on main results, which is subdivided into four subsections for ease of reading. The first of these concerns maximum principles for Frobenius and operator norms of matrix-valued functions (Theorems 3.1 and 3.3). Maximizing vectors (Definition 3.5) and an SVD type decomposition result for matrix-valued regular functions are brought out next (Theorem 3.8). We then discuss maximum principles involving singular values of matrix-valued functions (Theorem 3.12). Finite Blaschke products over $\mathbb{H}$ are introduced next (Definition 3.15). The last subsection concerns approximating a regular function $f:\mathbb{B}\rightarrow\overline{\mathbb{B}}$ that are continuous on $\partial\mathbb{B}$ by convex combinations of finite Blaschke products (Theorem 3.17). We end the paper by proving that an $n\times n$ matrix-valued regular function on the quaternionic unit ball $\mathbb{B}$, where each entry satisfies the above mentioned condition can be aproximated by matrix-valued function finite Blaschke products (Theorem 3.20). Wherever possible, we provide examples to illustrate the results obtained.
@@ 3.1. The Frobenius and Operator norm case
As pointed out in Theorem $2.12$ of [11], if we work with an arbitrary domain $\Omega$ in $\mathbb{H}$, then any regular function will have a power series expansion in each $\Sigma$-ball in $\Omega$ and in such a case, the proof of Theorem 3.3 can be modified appropriately.
@@ 3.2. Maximizing vectors and a decomposition theorem
Let $\Omega$ be a slice domain in $\mathbb{H}$ and $F:\Omega\rightarrow M_{n}(\mathbb{H})$ a regular matrix-valued function. If there exists a $q_{0}$ in $\Omega_{I}$ such that $\|F(q)\|\leq\|F(q_{0})\|$ for all $q\in\Omega$ and $x_{0}$ is the maximizing vector for $F(q_{0})$, then $F^{(k)}(q_{0})x_{0}=0$ for all $k\geq 1$.
@@
Let the assumptions stated above hold. Choose $R>0$ such that $\Sigma(q_{0},R)$ contained in $\Omega$ such that
@@
where $F_{n}\in M_{n}(\mathbb{H})$. Since $\mathbb{H}$ is a left Hilbert space, we have for any $x\in\mathbb{H}$,
@@
Let $\delta_{I}(q_{0},R)=\Sigma(q_{0},R)\cap L_{I}\subseteq\Sigma(q_{0},R)$ be the disk in $L_{I}$ of radius $R$ and centered at $q_{0}$. Then for any $q\in\delta_{I}(q_{0},R)$, choose $r<R$ such that $q=q_{0}+re^{it}$. We then have,
@@
By virtue of a regular power series representation of $F(q)$ , $\|F(q)\|\leq\|F(q_{0})\|=\|F_{0}\|$ for all $q\in\delta_{I}(q_{0},R)$ (see [11]); in particular, for the vector $x$ and $r<R$, it follows that
@@
Replacing $x$ by the maximizing vector $x_{0}$ of $F(q_{0})$, we get
@@
This implies that $F_{k}x_{0}=0$ for all $k\geq 1$ and hence $F^{(k)}(z_{0})x_{0}=0$ for all $k\geq 1$ and hence $F(q)x_{0}=F_{0}x_{0}=F(q_{0})x_{0}$ for all $q\in\delta_{I}(q_{0},R)$. We finally conclude from the identity principle (Theorem 2.8) that $F(q)x_{0}$ is constant on $\Omega$, as required. ∎
@@
It should be pointed out that we have assumed that the point $q_{0}\in\Omega_{I}$ (the point where the maximum is attained for $|F(q)|$) for some $I\in\mathbb{S}$ in the theorems proved in this section and the previous one. However, this is merely a technical issue and the proofs will go through verbatim, with appropriate modifications, if $q_{0}\in\Omega$. This is because for any $q\in\mathbb{H}$, we can choose $I,J\in\mathbb{S}$ such that $q\in L_{I}$ by taking $v=x_{1}i+x_{2}j+x_{3}k$ (the imaginary part of $q$) and by letting $I=\dfrac{v}{|v|}$ so that $q=x_{0}+|v|I\in L_{I}$.
@@ 3.4. Approximation of a quaternionic matrix-valued function
This is the last section of the paper. We prove as an application of Theorem 3.8, a Fisher type approximation theorem for matrix-valued regular functions. We begin this section by defining the Blaschke factor in the quaternion case (see [4] for details). Recall that the quaternionic Hardy space $H^{2}$ of the unit ball $\mathbb{B}$ is defined as the space of square summable regular power series:
$H^{2}:=\displaystyle\left\{f(q)=\sum_{n=0}^{\infty}q^{n}f_{n}:\|f\|^{2}_{H^{2}}:=\sum_{n=0}^{\infty}q^{n}|f_{n}|^{2}<\infty\right\}.$
Notice that for $\alpha\in\mathbb{B}$, the power series $k_{\alpha}(q)=\displaystyle\sum_{n=0}^{\infty}q^{n}\alpha^{n}$ is in $H^{2}$ and $\|k_{\alpha}\|^{2}_{H^{2}}=\displaystyle\frac{1}{1-|\alpha|^{2}}$.
@@
The power series $b_{\alpha}(q):=(q-\alpha)*k_{\bar{\alpha}}(q)$ is defined as the quaternionic analog of a Blaschke factor. The power series
$B(q):=b_{\alpha_{1}}(q)*b_{\alpha_{2}}(q)*\cdots*b_{\alpha_{n}}(q)\phi,\text{ where }|\phi|=1,\alpha_{i}\in\mathbb{B}\text{ for all }1\leq i\leq n$
is called a finite Blaschke product.
Note that this notion coincides with the usual definition of a Blaschke product over the field of complex numbers. We shall make use of the following approximation theorem attributed to Fisher (see Theorem $4.2.2$, [10]) in the following theorem.
@@
Each $f$ in the unit ball of the disc algebra $\mathcal{A}(\mathbb{D})$ can be uniformly approximated on $\overline{\mathbb{D}}$ by a convex combination of finite Blaschke products.
We are now in a position to prove a Fisher-type approximation theorem in the quaternion case.
Theorem 3.17.
Let $\mathbb{B}$ be the open unit ball in $\mathbb{H}$ and $f:\mathbb{B}\rightarrow\overline{\mathbb{B}}$ be a regular function on $\mathbb{B}$ such that $f$ is continuous on $\overline{\mathbb{B}}$. Then $f$ can be approximated in norm by a convex combination of finite Blaschke products.
@@
By the splitting lemma (Lemma 2.3), any regular function $f$ can be written as $f_{J}=g+hK$, where $J,K\in\mathbb{S},\ J\perp K$ and $g,h:\mathbb{B}_{J}\rightarrow\overline{\mathbb{B}_{J}}$ are holomorphic functions on $\mathbb{B}_{J}$ and continuous on $\overline{\mathbb{B}_{J}}$. This then implies that $g$ and $h$ are in the unit ball of the disc algebra $\mathcal{A}(\mathbb{D})$ (by identifying $\mathbb{B}_{J}$ with the unit disc $\mathbb{D}$). By Theorem 3.16 (see also [2]) for the complex case, $g$ and $h$ can be approximated by convex combinations of finite Blaschke products in field of complex numbers, say $\left\{B_{i}\right\}_{\{i=1\}}^{n}$ and $\{C_{j}\}_{\{j=1\}}^{m}$ such that $r_{1}(z)=\displaystyle\sum_{i=1}^{n}\lambda_{i}B_{i}(z)$ and $r_{2}(z)=\displaystyle\sum_{j=1}^{m}\mu_{j}C_{j}(z)$, where $\sum_{i=1}^{n}\lambda_{i}=\sum_{j=1}^{m}\mu_{j}=1$. Let us choose these functions so that for $\epsilon>0$,
$\|g(z)-r_{1}(z)\|<\frac{\epsilon}{8}\text{ and }\|h(z)-r_{2}(z)\|<\frac{\epsilon}{8},$
for all $z=x+yJ\in\mathbb{B}_{J}$. Since $K$ is a unimodular constant, each $D_{j}:=C_{j}K$ is a finite Blaschke product. Then define
$r(x+yJ):=\displaystyle\sum_{i=1}^{n}\chi_{i}B_{i}(x+yJ)+\sum_{j=1}^{m}\gamma_{j}D_{j}(x+yJ)+\frac{k}{2}(1)+\frac{k}{2}(-1),$
which is a convex combination of finite Blaschke products $\left\{B_{i}\right\}_{\{i=1\}}^{n}$ and $\left\{D_{j}:=C_{j}K\right\}_{\{j=1\}}^{m}$ in quaternions, where we choose $\chi_{i}$, $\gamma_{j}$ and $k$ satisfying the following conditions:
  • $\chi_{i}<\frac{\lambda_{i}}{2}$, $\gamma_{j}<\frac{\mu_{j}}{2}$, such that for given $\epsilon>0$, $|\sum_{i=1}^{n}(\lambda_{i}-\chi_{i})|<\epsilon$, $|\sum_{j=1}^{m}(\mu_{j}-\gamma_{j})|<\epsilon$, which is always possible, as $C_{j}$’s and $D_{j}$’s are finite Blaschke products and are bounded by $1$.
  • If $\sum_{i=1}^{n}\chi_{i}+\sum_{j=1}^{m}\gamma_{j}=1$ then take $k=0$; otherwise take $k=1-\displaystyle\sum_{i=1}^{n}\chi_{i}-\sum_{j=1}^{m}\gamma_{j}$.
We thus have $r(x+yJ):=\widetilde{r_{1}}(x+yJ)+\widetilde{r_{2}}(x+yJ)K$, with $\widetilde{r_{1}}(x+yJ)=\displaystyle\sum_{i=1}^{n}\chi_{i}B_{i}(x+yJ)$ and $\widetilde{r_{2}}(x+yJ)=\displaystyle\sum_{j=1}^{m}\gamma_{j}C_{j}(x+yJ)$, which is clearly a convex combination of finite Blaschke products. From what we have just done above, we see that for a given $\epsilon>0$,
$\|r_{1}(z)-\widetilde{r_{1}}(z)\|=\Big\|\displaystyle\sum_{i=1}^{n}\lambda_{i}B_{i}(z)-\sum_{i=1}^{n}\chi_{i}B_{i}(z)\Big\|\leq\Big|\sum_{i=1}^{n}(\lambda_{i}-\chi_{i})\Big|\Big\|B_{i}(z)\Big\|\leq\frac{\epsilon}{8}.$
By similar argument, $\|r_{2}(z)-\widetilde{r_{2}}(z)\|\leq\frac{\epsilon}{8}$. We can extend this function to the entire unit ball $\mathbb{B}$, which we again denote by $r(q)$, for $I\in\mathbb{S}$ (see Proposition $1.3.2$ of [9] for details).; that is,
$ext(r(q))=r(x+yI)=\frac{1}{2}\big[r(x+yJ)+r(x-yJ)+IJ\left[r(x-yJ)-r(x+yJ)\right]\big].$
@@
$|f(q)-r(q)|$ $=\Bigg|\frac{1}{2}[[f(x+yJ)+f(x-yJ)]+IJ\left[f(x-yJ)-f(x+yJ)\right]$
@@
$\|f(q)-r(q)\|$ $=\Bigg\|\frac{1}{2}\Bigg[(1-IJ)\Big[g(x+yJ)-\widetilde{r_{1}}(x+yJ)+(h(x+yJ)-\widetilde{r_{2}}(x+yJ))K\Big]$
$+(1+IJ)\Big[g(x-yJ)-\widetilde{r_{1}}(x-yJ)+(h(x-yJ)-\widetilde{r_{2}}(x-yJ))K\Big]\Bigg]\Bigg\|$
$\leq\Big\|g(x+yJ)-\widetilde{r_{1}}(x+yJ)\Big\|+\Big\|h(x+yJ)-\widetilde{r_{2}}(x+yJ)\Big\|$
$+\Big\|g(x-yJ)-\widetilde{r_{1}}(x-yJ)\Big\|+\Big\|h(x-yJ)-\widetilde{r_{2}}(x-yJ)\Big\|$
$\leq\Big\|g(x+yJ)-r_{1}(x+yJ)\Big\|+\Big\|r_{1}(x+yJ)-\widetilde{r_{1}}(x+yJ)\Big\|$
$+\Big\|h(x+yJ)-r_{2}(x+yJ)\Big\|+\Big\|r_{2}(x+yJ)-\widetilde{r_{2}}(x+yJ)\Big\|$
$+\Big\|g(x-yJ)-r_{1}(x-yJ)\Big\|+\Big\|r_{1}(x-yJ)-\widetilde{r_{1}}(x-yJ)\Big\|$
$+\Big\|h(x-yJ)-r_{2}(x-yJ)\Big\|+\Big\|r_{2}(x-yJ)-\widetilde{r_{2}}(x-yJ)\Big\|$
@@
We thus conclude that $f$ can be approximated by a convex combination of finite Blaschke products over $\mathbb{H}$, as desired. ∎
@@
We wish to point out that a classical approximation result due to Caratheodory says that any element in the unit ball of $H^{\infty}$ can be uniformly approximated by sequence of finite Blaschke products on compact subsets of the unit ball of $\mathbb{C}$ (see Theorem $4.1.1$, [10]). One can then prove the following theorem. We skip the proof.
@@
Let $\mathbb{B}$ be the open unit ball in $\mathbb{H}$ and $f:\mathbb{B}\rightarrow\overline{\mathbb{B}}$ be a regular function on $\mathbb{B}$. Then $f$ can be approximated by a sequence of finite Blaschke products on every compact subset of the unit ball.
We conclude this section by extending Theorem 3.17 to matrix-valued regular functions.
Theorem 3.20.
Let $F:\mathbb{B}\rightarrow M_{n}(\mathbb{H})$ be a regular function such that $\|F(q)\|\leq 1$ with all entries being continuous on $\overline{\mathbb{B}}$. Then $F$ can be approximated by a convex combination of $M_{n}(\mathbb{H})$-valued finite Blaschke products.
@@
We will prove this by induction on the order of matrix $n$. Since $\|F(q)\|\leq 1$, Theorem 3.8 tells us that there exist $n\times n$ constant unitary matrices $U$ and $V$ and a regular function $G:\mathbb{B}\rightarrow M_{n-1}(\mathbb{H})\cong\mathbb{H}$ such that
$F(q)=U\begin{bmatrix}1&0\\ 0&G(q)\end{bmatrix}V^{*}.$
Note that $\|G(q)\|\leq 1$; moreover the entries of $G$ are also continuous on $\overline{\mathbb{B}}$. When $n=2$, Theorem 3.17 implies that there is an approximation of $G$ by convex combination of finite Blaschke products over $\mathbb{H}$. This then in turn implies that $F$ can also be approximated by a convex combination of matrix-valued finite Blaschke products (notice that $U$ and $V$ are constant unitary matrices of norm $1$). Suppose the result is true for n=k1n=k-1 and let $F(q)$ be a $k\times k$ matrix-valued function satisfying the assumptions of the theorem. Decompose $F(q)$ as above. Then by the decomposition theorem, $G(q)$ which is $(k-1)\times(k-1)$ matrix-valued regular function with norm bounded by $1$, can be approximated by a convex combination of matrix-valued finite Blaschke products (induction argument). Then $F$ can also be approximated by a convex combination of matrix-valued finite Blaschke products. ∎
We end by making some remarks.
Remark 3.21.
  • Notice that an $n\times n$ matrix whose entries are themselves finite Blaschke products is not a matrix-valued finite Blaschke product. However, this issue does not arise in the above theorem, as we first reduce the function $F$ to a diagonal form using constant unitary matrices (having norm one) and only approximate the last term in the decomposition by finite Blaschke products, thereby giving a matrix-valued finite Blaschke product.
  • A suitable version of Theorem 3.20 can be proved when the function $F$ is merely a norm one matrix-valued regular function, where convergence happens over every compact subset of the unit ball.
  • Regular functions that map the quaternionic unit ball to itself has been extensively studied by Alpay et al in [1] in the context of interpolation problems.